OpenStudy (anonymous):

The table below shows what 300 Boxer students ate for lunch on Friday. What is the ratio of the number of students who ate pizza to the number of students who did not eat pizza? SoupPizzaSalad Hotdog 60 120 30 90 A.3/2 B.1/4 C.2/3 D.2/5 E.5/2 F.1/6

OpenStudy (texaschic101):

120/180 = 2/3 (I think) Because 120 ate pizza and 180 didn't.

OpenStudy (anonymous):

120/180 divided by 2/2 = 60/90 then divide that by 3/3= 20/30 then divide by 10/10 and you get 2/3 texaschic101 was write i just wanted to show you how to get that answer